330 views
LHC beam instrumentation status =============================== ###### tags: `LHC` 9 April 2021 https://indico.cern.ch/event/1025164/ ### Actions: Enrico to check with OP about access in IP4. Enrico and Belen to organize a TB about the on demand PM events. ## LHC schedule The machine will be ready for beam in week 38. Training of sector 67 being the latest task. RF conditioning between weeks 29 and 32 included, LLRF tuning in weeks 33 and 34. Access inside the RF zone (inside the yellow grids) not possible during this time. Access may be possible between weeks 35 and 38 included, to be discussed with OP. ### Pilot Beam Test Scheduled for weeks 39 and 40 (start on September 27). Could be posponed by up to two weeks, but impact on retart to be evaluated. 3 days of checkout, this is only for the systems that require the beam permit loop closed. All other systems will be tested before. 10 days of operation with beam. Up to 3 nominal bunches at injection, one probe bunch at FT. Collisions at injection in stable beams mode. ### Run 3 start Experimental cavers closed by February 1st, checkout and beam commissioning right after. After the pilot test there will be a *Maintenance period*, basically the same as a YETS, therefore activities in that period should be planned in the same way as for a YETS. ## BI Status Most systems are on track, still some work to be done, but either minor/standard things or validation work. The systems requiring more work are the BLM, the BTV, the BRAN and the BWS. ### BLM * Repair optical links (~30) * Install/verify FIP boxes * Finalize PM, XPOC and injection-inhibit functionalities * Modify LAYOUT to upload data to LSA * Modify InforEAM and LAYOUT to remove MTF * Update thresholds (some work still needed on the tools) ### BTV FESA class same as for injectors for most BTV systems (ready). FESA class for BTVDD need to be ported/updated. ### Orbit and tune feedback Major overhaul of the controller (now single FESA class BFCLHC). System is complete, but needs to be validated (with OP). ### BRAN * BRAN-A (berkely IC) only needs to be restarted (FESA class porting status?), gas, safety relief valves etc. * BRAN-C OK (FESA class porting status?). * BRAN-D, this is a new design that replaces 2 BRAN-A ionization chambers. Still assembling the detectors, some issues encountered. Acquisition FW not started yet. For the pilot beam test the BRAN-A are not useable, would be nice to have the BRAN-D. Eventually the detectors can be read out using the electronics and FESA of the BRAN-C (only for the test beam). ### BWS Two (H) scanners have been modified and will be equipped with a new control and acquisition system (same as LIU scanners). The 6 operational scanners will remain as in run 2. The FESA class of the operational devices has to be adapted (same as for the SPS linear scanners). The modified systems still require a lot of work: install the electronics, provide a control system etc. ## Next steps There will be a presentation at the LBOC on April 20 about the status of the LHC beam instrumentation. The questions to answer are: 1) HW changes during LS2 2) SW changes during LS2 3) Status of the HW installation 4) Status of the SW 5) When can OP start testing the system 6) Is there an established procedure for testing the system? Please send the slides. There will be another presentation at the LMC on May 12 (basically the same). The next meeting will be around May 21, exact date to be confirmed.